OpenStudy (anonymous):

Write the equation of the line that passes through (1, 3) and (4, 4) in standard form

Directrix (directrix):

1*x + (-3) y = 8

hero (hero):

\[\text{1. Find the slope, m, 'tween the points} \space m = \frac{y_2 - y_1}{x_2 - x_1}\]

hero (hero):

Then use \[(y - y_1) = m(x - x_1)\]

Directrix (directrix):

slope = 1/3 y-intercept 8/3 y = (1/3)x + 8/3 3y = x + 8

Directrix (directrix):

Standard Form: the standard form of a line is in the form Ax + By = C where A is a positive integer, and B, and C are integers.